OpenStudy (anonymous):

Colin has a collection of 120 comic books. Fifty-four of the comic books were damaged in a flood. What percent of the comic books were damaged?

OpenStudy (butterflydreamer):

For this question you'd answer it by doing this: 54/120 = ? Then mutiply that number by 100 to get your percentage :)

OpenStudy (anonymous):

54/100 =.54 *100 is 54%

OpenStudy (butterflydreamer):

no the total books are 120... so 54/120 and then multiply that by 100 to get the percentage

OpenStudy (butterflydreamer):

54/120 = 0.45 Then multiple that by 100 0.45 * 100 = 45 % NOT 54 %

OpenStudy (butterflydreamer):

@AidanWithanAnota

OpenStudy (anonymous):

Hi